# REIC: RAG-Enhanced Intent Classification at Scale

### v1

@@ Abstract.
Intent Classification, RAG, ICL, text classification
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Classification and regression treesccs: Computing methodologies Natural language processing
1. Introduction
Customer service (Cui et al., 2017; Yu et al., 2021a; Nguyen et al., 2020; Qi et al., 2021; Parikh et al., 2023; Zhou et al., 2023; Chen et al., 2024) is critical for modern e-commerce but also one of the most resource-intensive departments. Different agents, either human or model, are trained to handle specific types of customer issues, making precise intent classification, particularly at the issue level, crucial for efficient routing. High issue-oriented intent accuracy ensures that customers are connected with the most suitable agents, reducing unnecessary transfers and lowering handling times. This optimization not only enhances customer satisfaction but also cuts operational costs by streamlining interactions and improving overall service efficiency. For model-based automatic resolvers in chatbot agentic systems (Gupta et al., 2024; Chen et al., 2023), the ability to precisely identify user intent is essential for delivering contextually appropriate and solution-oriented responses.
As companies expand their product lines, intent classification faces two key challenges. First, the number of customer intents grows over time, requiring models to adapt to new intents quickly. Second, intent taxonomies can vary across product lines, making it difficult to maintain a unified classification system. For example we organize products into different verticals in e-commerce: with third-party products, customer usually inquire about physical retail orders or consumer accounts, and intents are categorized into three levels from coarse to fine-grained. In contrast, first-party products require more customized customer services due to our proprietary device and digital product offerings. For instance, a customer seeking device troubleshooting may interact with an agent who can access real-time diagnostic information and perform specific troubleshooting steps on the user’s behalf. This necessitates a broader set of intent categories to accommodate diverse customer needs, as illustrated in Figure 1. This heterogeneity complicates intent classification, demanding scalable and flexible approaches to ensure accurate routing and efficient customer service. In this work we demonstrate our method using two verticals but it can be easily adapted to more.
@@ 1. Introduction
Figure 1. We present the heterogeneous intent structure with representative examples, illustrating the intent label hierarchy in each vertical.
2. Related Work
2.1. Intent classification
Early work on intent classification for dialogues often relied on bag-of-words or recurrent models. For example, Schuurmans and Frasincar (2019) evaluated various classifiers on a multi-domain intent dataset and found that a simple SVM with hierarchical label taxonomy outperformed deeper LSTM models. With the advance of transformer architectures, researchers began to leverage self-attention and multi-task learning for intent understanding. Ahmadvand et al. (2020) introduced a joint intent mapping model that simultaneously classifies high-level intent and maps queries to fine-grained product categories. Wang et al. (2021) employed a slowly updated text encoder and global/local memory networks to mitigate catastrophic forgetting and parameter explosion for large-scale intent detection task. Recent work has pushed toward using large pre-trained models and retrieval-based prompting to enable cross-domain and zero/few-shot intent classification. Liu et al. (2024) proposed a framework which integrates a fine-tuned XLM-based intent classifier with an LLM to essentially treat multi-turn intent understanding as a zero-shot task. Yu et al. (2021b) also explored retrieval-based methods for intent classification and slot filling tasks in few-shot settings. Our work adopts similar in-context learning (ICL) setup while focusing on handling large-scale multi-domain intent classification task from industry level applications.
2.2. In-context Learning
The performance of LLM has been significantly enhanced in few-shot and zero-shot NLP tasks through ICL. Recent ICL research focus on how to effectively identify and interpret retrieved context. Guu et al. (2020) first showed how to pre-train masked language models with a knowledge retriever in an unsupervised manner. Karpukhin et al. (2020) proposed a training pipeline in which retrieval is implemented using dense representations alone and embeddings are learned from a small number of questions and passages with a dual-encoder. Ram et al. (2023) considered simple alternatives to only prepend retrieved grounding documents to the input, instead of modifying the LLM architecture to incorporate external information. Similar approaches have proven particularly effective in the application of RAG on dialogue systems (Shuster et al., 2022b, a), specifically goal-oriented and domain-specific dialogs from customer service scenarios (Zhuang et al., 2021, 2023; Senel et al., 2024). In our work, we utilize ICL in both LLM fine-tuning stage for data generation and at inference-time with an intent candidate retriever.
Figure 2. The proposed REIC method from customer query to routing intent leveraging on vector retrieval and probability calculation.
3. Preliminary
@@
Flat classification assumes independence among intent labels, which is often unrealistic in practice. Two key challenges arise at industry scale: 1) Scalability—a large number of labels makes flat classification inefficient and hard to maintain; 2) Label correlation—related intents (e.g., “Order Issue” vs. “Track Order”) are treated independently, ignoring useful hierarchical structure. To address this, we adopt hierarchical intent classification, where queries are classified progressively from general to specific intents. This improves both efficiency and accuracy.
Note that for a more generalized setting, intents from different verticals and domains may have entirely different hierarchy and ontology. In Figure 1, we demonstrated some examples of intent hierarchy in our application, which involves customer service query intent detection with two business verticals: Third-Party or 3P business (customer contacting about third-party physical retail orders or consumer accounts) and First-Party or 1P business (customer contacting about first-party digital or device issues). If we use the traditional single-head flattened intent labels, the total intent ontology set size would be at $10^{3}$ level which create major challenges for accurate intent classification. By creating hierarchical intent ontology across different business verticals, each classification head only needs to handle less than 50 intents that are more manageable for language models. In the following sections of this paper, we utilize this intent ontology setup for experiments and comparisons.
4. Method
Adapting Large Language Models (LLMs) to domain-specific intent detection is challenging due to specialized terminology, organizational language, and unique customer scenarios. To address these limitations and enhance the accuracy of customer intent identification, we introduce a novel approach REIC for RAG-Enhanced Intent Classification. REIC consists of three components: index construction, candidate retrieval, and intent probability calculation (as illustrated in Figure 2).
@@ 4. Method
(2) $sim(q,q_{i})=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{i}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{i}\|}$
@@
Figure 3. Constrained decoding for probability calculation.
@@
(5) $\mathcal{M}(\mathcal{P},q,E)_{t_{j}}=\exp{(\sum_{t_{j}}\text{LogSoftmax}(\mathcal{L}_{t_{j}})/s_{j})}$
@@
5. Experimental Setup
5.1. Datasets
We applied data anonymization described in Appendix A and the final dataset contains 52,499 training samples with 35,041 1P Business queries and 17,458 3P Business queries. The test set consists of 3,647 1P Business queries and 1,717 3P Business queries respectively using random sampling. All of the data samples have incorporated retrieved intent candidates from the retriever. We also performed dataset cleaning in the training set to make sure the true intent is contained in the retrieved list. During inference, we use the actual noisy retrieved list which also relies on the capability of the embedding model.
Table 1. Intent detection confusion matrix on different business with different methods
@@ 5.1. Datasets
5.2. Compared Methods
@@
  • RoBERTa: We fine-tune RoBERTa-base (Liu et al., 2019) with multiple classification heads. This adaptation allowed the model to simultaneously categorize utterances across multiple dimensions.
@@
5.3. Implementation Details
The LLM component of our REIC approach utilizes a fine-tuned model from Mistral-7B-Instruct-v0.2333https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2. The retrieval index is implemented using FAISS (Douze et al., 2024) for efficient similarity search in high-dimensional space. We experimented with four different retrievers: BM25 (Robertson et al., 1995), MPNet (Song et al., 2020), Contriever-MS MACRO (Izacard et al., 2022) and ColBERT-v2 (Santhanam et al., 2022). More implementation details are described in Appendix D.
6. Results
6.1. Intent Detection Ability
To assess REIC’s effectiveness in intent detection, we compared it against baseline methods (see §5.2), with results shown in Table 1. The experiments, covering the 3P Business and 1P Business verticals, evaluated Precision, Recall, and F1-score.
The results indicate that REIC outperforms standard fine-tuning and prompting-based methods. While fine-tuned models like RoBERTa perform reasonably well, they require extensive retraining when new intents emerge. We hypothesize that the limited performance of the Mistral Classification model stems from its nature as a decoder-only architecture, which may be less effective in extracting the semantic meaning of input query. Prompting-based approaches (Claude Zero-shot and Few-shot) generally underperformed, with Claude Few-shot achieving a maximum F1-score of 0.329 overall. The Claude + RAG method improved performance compared to standalone prompting but remained inferior to our approach by 26.7%.
These observations confirm that the integration of RAG and fine-tuned LLM enables greater flexibility, improved precision, and higher recall rates, making it well-suited for handling diverse and evolving intent spaces in different applications. We further evaluate the impact of different retrievers and the size of retrieval candidates in Appendix B and Appendix C.
6.2. Robustness on Unseen Intents
@@
As seen in the table, Claude Zero-shot performs poorly with 0.17 accuracy, while Claude Few-shot improves to 0.308. RAG-based methods, particularly Claude + RAG, outperform Claude Few-shot with 0.389 accuracy, demonstrating the advantage of dynamic retrieval. REIC, though slightly lower than Claude + RAG, still outperforms Claude Few-shot, underscoring its strong performance on both in-domain and out-of-domain intents. Overall, REIC proves robust and adaptable to unseen domains.
Table 2. Out-of-domain intent detection accuracy
@@
7. Conclusion
@@
Appendix
Appendix A Data Anonymization
Due to business considerations, we are not permitted to share the results using the original customer data. As a result, we manually anonymized both the labels and transcripts to ensure no personal information is included. Additionally, specific product and service names were denonymized to prevent the identification of the company from the transcript or label descriptions. Despite these modifications, the conclusions drawn from our experiments remain valid.
Appendix B Impact of Retrievers
In order to evaluate the impact of retrievers on the final performance, we experimented four different retrievers in REIC including one sparse retrieval method and three dense retrievers, details in Appendix D. The intent detection accuracy across different business verticals using these retrievers is presented in Table 3.
Our findings show that retriever selection significantly impacts intent classification. Although BM25 is a strong baseline, dense retrievers like MPNet consistently outperform it. This highlights the value of high-quality embeddings and extensive fine-tuning on large datasets, which is why we have chosen MPNet as our final retriever in REIC. We also conducted experiments to study the impact of retrieval candidate size in Appendix C.
Table 3. Intent detection accuracy on different business verticals using different retrievers in REIC
Retriever 3P Business 1P Business Overall
BM25 0.521 0.537 0.532
Contriever 0.450 0.461 0.457
ColBERTv2 0.503 0.560 0.542
MPNet 0.545 0.573 0.564
Appendix C Impact of Retrieval Candidate Size
Figure 4. The accuracy and latency when using different retrieval top-k values.
Appendix D Implementation Details
To fine-tune our REIC LLM, we apply 8 NVIDIA-A100 40GB GPUs with 96 vCPUs to conduct PEFT (Mangrulkar et al., 2022) training with LoRA adapters (Hu et al., 2022). We choose a set of LoRA parameters with a rank of 8, an alpha value of 16, and a dropout rate of 0.1. The training batch size is set to 8 per GPU with a learning rate of $2e^{-5}$. We train the model using Cross Entropy Loss for 3 epochs which takes around 3 hours on the instance.
  • BM25 (Robertson et al., 1995) is a widely used traditional sparse retrieval method. Although it is unsupervised, it consistently demonstrates strong performance across a variety of benchmarks(Thakur et al., 2021).
  • MPNet [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 (Song et al., 2020) is a sentence embedding model fine-tuned on one billion sentence pairs using a contrastive learning objective.
  • Contriever-MS MACRO (Izacard et al., 2022) is an unsupervised dense retriever pre-trained with contrastive learning and fine-tuned on MS MARCO (Nguyen et al., 2016).
  • ColBERT-v2 (Santhanam et al., 2022) is a late-interaction retriever that combines denoised supervision and residual compression to improve retrieval quality and reduce space footprint.

### v2

@@ Abstract
REIC: RAG-Enhanced Intent Classification at Scale
Ziji Zhang, Michael Yang, Zhiyu Chen, Yingying Zhuang, Shu-Ting Pi, Qun Liu, Rajashekar Maragoud, Vy Nguyen, Anurag Beniwal Amazon.com Inc, Seattle, USA {czhangzi,abyang,zhiyuche,yyzhuang,shutingp}@amazon.com {qunliu,maragoud,nguynvy,beanurag}@amazon.com
1 Introduction
Customer service cui2017superagent; chen-etal-2024-identifying; nguyen-etal-2020-dynamic; qi-etal-2021-benchmarking; chen-etal-2023-generate; zhou-etal-2023-towards-open; pi2024topology is critical for modern e-commerce but also one of the most resource-intensive departments. Different agents, either human or model, are trained to handle specific types of customer issues, making precise intent classification, particularly at the issue level, crucial for efficient routing. High issue-oriented intent accuracy ensures that customers are connected with the most suitable agents, reducing unnecessary transfers and lowering handling times. This optimization not only enhances customer satisfaction but also cuts operational costs by streamlining interactions and improving overall service efficiency. For model-based automatic resolvers in chatbot agentic systems gupta2024dard, the ability to precisely identify user intent is essential for delivering contextually appropriate and solution-oriented responses.
As companies expand their product lines, intent classification faces two key challenges. First, the number of customer intents grows over time, requiring models to adapt to new intents quickly. Second, intent taxonomies can vary across product lines, making it difficult to maintain a unified classification systempi2024universal. For example we organize products into different verticals in e-commerce: with third-party products, customer usually inquire about physical retail orders or consumer accounts, and intents are categorized into three levels from coarse to fine-grained. In contrast, first-party products require more customized customer services due to our proprietary device and digital product offerings. For instance, a customer seeking device troubleshooting may interact with an agent who can access real-time diagnostic information and perform specific troubleshooting steps on the user’s behalf. This necessitates a broader set of intent categories to accommodate diverse customer needs, as illustrated in Figure 1. This heterogeneity complicates intent classification, demanding scalable and flexible approaches to ensure accurate routing and efficient customer service. In this work we demonstrate our method using two verticals but it can be easily adapted to more.
@@ 1 Introduction
Figure 1: We present the heterogeneous intent structure with representative examples, illustrating the intent label hierarchy in each vertical.
2 Related Work
2.1 Intent classification
Early work on intent classification for dialogues often relied on bag-of-words or recurrent models. For example, schuurmans2019intent evaluated various classifiers on a multi-domain intent dataset and found that a simple SVM with hierarchical label taxonomy outperformed deeper LSTM models. With the advance of transformer architectures, researchers began to leverage self-attention and multi-task learning for intent understanding. ahmadvand2020jointmap introduced a joint intent mapping model that simultaneously classifies high-level intent and maps queries to fine-grained product categories. wang2021mell employed a slowly updated text encoder and global/local memory networks to mitigate catastrophic forgetting and parameter explosion for large-scale intent detection task. Recent work has pushed toward using large pre-trained models and retrieval-based prompting to enable cross-domain and zero/few-shot intent classification. liu2024lara proposed a framework which integrates a fine-tuned XLM-based intent classifier with an LLM to essentially treat multi-turn intent understanding as a zero-shot task. yu2021few also explored retrieval-based methods for intent classification and slot filling tasks in few-shot settings. Our work adopts similar in-context learning (ICL) setup while focusing on handling large-scale multi-domain intent classification task from industry level applications.
2.2 In-context Learning
The performance of LLM has been significantly enhanced in few-shot and zero-shot NLP tasks through ICL. Recent ICL research focus on how to effectively identify and interpret retrieved context. guu2020realmretrievalaugmentedlanguagemodel first showed how to pre-train masked language models with a knowledge retriever in an unsupervised manner. karpukhin2020densepassageretrievalopendomain proposed a training pipeline in which retrieval is implemented using dense representations alone and embeddings are learned from a small number of questions and passages with a dual-encoder. ram2023incontextretrievalaugmentedlanguagemodels considered simple alternatives to only prepend retrieved grounding documents to the input, instead of modifying the LLM architecture to incorporate external information. Similar approaches have proven particularly effective in the application of RAG on dialogue systems shuster2022blenderbot3deployedconversational; shuster-etal-2022-language, specifically goal-oriented and domain-specific dialogs from customer service scenarios zhuang-etal-2021-weakly; 10.1145/3543873.3587680. In our work, we utilize ICL in both LLM fine-tuning stage for data generation and at inference-time with an intent candidate retriever.
Figure 2: The proposed REIC method from customer query to routing intent leveraging on vector retrieval and probability calculation.
3 Preliminary
@@
However, intent labels are often not mutually independent; correlations between intents can exist, making a flat classification structure suboptimal. This leads to two major challenges in large-scale, industry-level intent classification: 1) scalability: a large number of intent labels $k$ can make flat classification computationally expensive and difficult to scale, especially as the set of intents grows; 2) label correlation: related intents, such as "Order Issue" and its sub-intents “Track Order” or “Cancel Order”, are treated independently in flat classification, ignoring their hierarchical relationships and increasing the risk of misclassification.
To address these issues, hierarchical intent classification is preferred. In this approach, a query $q$ is classified progressively from general categories to specific sub-intents, enhancing both efficiency and accuracy for industry-scale applications.
Note that for a more generalized setting, intents from different verticals and domains may have entirely different hierarchy and ontology. In Figure 1, we demonstrated some examples of intent hierarchy in our application, which involves customer service query intent detection with two business verticals: Third-Party or 3P business (customer contacting about third-party physical retail orders or consumer accounts) and First-Party or 1P business (customer contacting about first-party digital or device issues). Both verticals span a diverse range of product types, reflecting the broad scope of customer inquiries handled by our system. If we use the traditional single-head flattened intent labels, the total intent ontology set size would be at $10^{3}$ level which create major challenges for accurate intent classification. By creating hierarchical intent ontology across different business verticals, each classification head only needs to handle less than 50 intents that are more manageable for language models. In the following sections of this paper, we utilize this intent ontology setup for experiments and comparisons.
4 Method
LLMs has revolutionized the landscape of customer engagement, particularly in the domain of intent detection systems. While these models demonstrate remarkable capabilities in language comprehension and knowledge representation, their adaptation to domain-specific contexts presents notable challenges. Specifically, the integration of industry-specific terminology, organizational nomenclature, and distinctive customer service scenarios necessitates fine-tuning and customization of these models. To address these limitations and enhance the accuracy of customer intent identification, we introduce a novel approach REIC for RAG-Enhanced Intent Classification.
REIC aims to bridge the gap between the generalized capabilities of LLMs and the specialized requirements of diverse business environments. By leveraging RAG, our approach seeks to enhance the precision and relevance of intent detection, thereby facilitating more nuanced and context-aware customer interactions across various industry-specific scenarios. The method consists of three main components: index construction, candidate retrieval, and intent probability calculation (Figure 2).
@@ 4 Method
$sim(q,q_{i})=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{i}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{i}\|}$ (2)
@@
Figure 3: Constrained decoding for probability calculation.
@@
$\mathcal{M}(\mathcal{P},q,E)_{t_{j}}=\exp{(\sum_{t_{j}}\text{LogSoftmax}(\mathcal{L}_{t_{j}})/s_{j})}$ (5)
@@
5 Experimental Setup
5.1 Datasets
Due to business considerations, we are not permitted to share the results using the original customer data. As a result, we manually anonymized both the labels and transcripts to ensure no personal information is included. Additionally, specific product and service names were denonymized to prevent the identification of the company from the transcript or label descriptions. Despite these modifications, the conclusions drawn from our experiments remain valid. The final dataset contains 52,499 training samples with 35,041 1P Business queries and 17,458 3P Business queries. The test set consists of 3,647 1P Business queries and 1,717 3P Business queries respectively using random sampling. All of the data samples have incorporated retrieved intent candidates from the retriever. We also performed dataset cleaning in the training set to make sure the true intent is contained in the retrieved list. During inference, we use the actual noisy retrieved list which also relies on the capability of the embedding model.
@@ 5.1 Datasets
Table 1: Intent detection confusion matrix on different business with different methods
5.2 Compared Methods
@@
  • RoBERTa: We fine-tune RoBERTa-base liu2019roberta with multiple classification heads. This adaptation allowed the model to simultaneously categorize utterances across multiple dimensions.
@@
5.3 Implementation Details
The LLM component of our REIC approach utilizes a fine-tuned model from Mistral-7B-Instruct-v0.2333https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2. We applied 8 NVIDIA-A100 40GB GPUs with 96 vCPUs to conduct PEFT peft training with LoRA adapters hu2022lora. We choose a set of LoRA parameters with a rank of 8, an alpha value of 16, and a dropout rate of 0.1. The training batch size is set to 8 per GPU with a learning rate of $2e^{-5}$. We train the model using Cross Entropy Loss for 3 epochs which takes around 3 hours on the instance.
  • BM25 robertson1995okapi is a widely used traditional sparse retrieval method. Although it is unsupervised, it consistently demonstrates strong performance across a variety of benchmarksthakur2021beir.
  • MPNet [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 song2020mpnet is a sentence embedding model fine-tuned on one billion sentence pairs using a contrastive learning objective.
  • Contriever-MS MACRO izacard2022unsupervised is an unsupervised dense retriever pre-trained with contrastive learning and fine-tuned on MS MARCO nguyen2016ms.
  • ColBERT-v2 santhanam-etal-2022-colbertv2 is a late-interaction retriever that combines denoised supervision and residual compression to improve retrieval quality and reduce space footprint.
6 Results
6.1 Intent Detection Ability
To evaluate the effectiveness of our REIC method in intent detection, we conducted experiments comparing it against several baseline methods described in §5.2. Our results, presented in Table 1, illustrate performance across two business verticals (3P Business and 1P Business) and an overall aggregate assessment based on Precision, Recall, and F1-score.
The results indicate that our REIC offers significant advantages in intent detection over standard fine-tuning or prompting-based methods. While fine-tuned models like RoBERTa perform reasonably well, they require extensive retraining when new intents emerge. We hypothesize that the limited performance of the Mistral Classification model stems from its nature as a decoder-only architecture, which may be less effective in extracting the semantic meaning of input query. Additionally, since it is not pretrained for classification tasks, incorporating a classification head during fine-tuning is unlikely to yield optimal results. Prompting-based approaches (Claude Zero-shot and Few-shot) generally underperformed, with Claude Few-shot achieving a maximum F1-score of 0.329 overall. The Claude + RAG method improved performance compared to standalone prompting but remained inferior to our approach by 26.7%.
These observations confirm that the integration of RAG and fine-tuned LLM enables greater flexibility, improved precision, and higher recall rates, making it well-suited for handling diverse and evolving intent spaces in different applications.
6.2 Impact of Retrievers
In order to evaluate the impact of retrievers on the final performance, we experimented four different retrievers in REIC including one sparse retrieval method and three dense retrievers, details in §5.3. The intent detection accuracy across different business verticals using these retrievers is presented in Table 2.
Our findings show that retriever selection significantly impacts intent classification. Although BM25 is a strong baseline, dense retrievers like MPNet consistently outperform it. This highlights the value of high-quality embeddings and extensive fine-tuning on large datasets, which is why we have chosen MPNet as our final retriever in REIC.
Retriever 3P Business 1P Business Overall
BM25 0.521 0.537 0.532
Contriever 0.450 0.461 0.457
ColBERTv2 0.503 0.560 0.542
MPNet 0.545 0.573 0.564
Table 2: Intent detection accuracy on different business verticals using different retrievers in REIC
7 Impact of Retrieval Candidate Size
Figure 4: The accuracy and latency when using different retrieval top-k values.
7.1 Robustness on Unseen Intents
@@
Claude Zero-shot performs the worst, with an accuracy of 0.17 on the 1P Business vertical. In contrast, Claude Few-shot shows improvement, achieving an accuracy of 0.308 on the 1P Business vertical. This demonstrates that providing a few examples significantly enhances the model’s ability to generalize.
Notably, the RAG-based methods, particularly Claude + RAG, significantly outperform Claude Few-shot, achieving 0.389 on the 1P Business vertical. This demonstrates the advantage of our RAG-based strategy in handling unseen intents, as it dynamically retrieves the most relevant examples to enhance predictions, surpassing static few-shot examples. Similarly, REIC, although slightly lower than Claude + RAG, still performs strongly compared to Claude Few-shot, highlighting the model’s effectiveness on both in-domain and out-of-domain intents. Overall, REIC excels in-domain, and its performance on the 1P Business vertical remains competitive with Claude Few-shot, underscoring the robustness and adaptability of our REIC approach for unseen domains.
@@
Table 3: Out-of-domain intent detection accuracy
7.2 Online Deployment Performance
Following the deployment of our REIC in an internal system, we observed significant improvements in online performance, particularly in intent detection accuracy. Compared to the previously deployed system, which relies on two separate models using traditional fine-tuning approaches for different business verticals, our REIC method simplified the system using just one consolidated model, which reduced misclassifications and improved resolution routing. To measure the success rate, we include a confirmation question for the customer to verify whether the predicted intent or issue is correct. Our REIC leads to a 3.38% absolute improvement for customer positive response rate. These improvements validate the efficacy of REIC for real-world deployment, offering both accuracy gains and operational efficiency in intent classification.
8 Conclusion
@@
9 Limitations
While REIC demonstrates strong performance in both in-domain and out-of-domain intent classification, it has a few limitations. Its accuracy remains closely tied to the quality of the retriever: if the correct intent is not among the retrieved candidates, the model cannot recover, underscoring the need for more robust retrieval methods or fallback mechanisms. In addition, although REIC allows for dynamic updates without retraining, it still relies on a fixed number of retrieved candidates, creating a trade-off between accuracy and latency that may hinder its deployment in real-time applications. Future work could address these challenges by developing adaptive retrieval strategies or introducing confidence-based mechanisms to dynamically adjust the candidate pool.